# Math-LLaVA: Bootstrapping Mathematical Reasoning for Multimodal Large Language Models

### v2

@@ Abstract
Large language models (LLMs) have demonstrated impressive reasoning capabilities, particularly in textual mathematical problem-solving. However, existing open-source image instruction fine-tuning datasets, containing limited question-answer pairs per image, do not fully exploit visual information to enhance the multimodal mathematical reasoning capabilities of Multimodal LLMs (MLLMs). To bridge this gap, we address the lack of high-quality, diverse multimodal mathematical datasets by collecting 40K high-quality images with question-answer pairs from 24 existing datasets and synthesizing 320K new pairs, creating the MathV360K dataset, which enhances both the breadth and depth of multimodal mathematical questions. We introduce Math-LLaVA, a LLaVA-1.5-based model fine-tuned with MathV360K. This novel approach significantly improves the multimodal mathematical reasoning capabilities of LLaVA-1.5, achieving a 19-point increase and comparable performance to GPT-4V on MathVista’s minitest split. Furthermore, Math-LLaVA demonstrates enhanced generalizability, showing substantial improvements on the MMMU benchmark. Our research highlights the importance of dataset diversity and synthesis in advancing MLLMs’ mathematical reasoning abilities. The code and data are available at: https://github.com/HZQ950419/Math-LLaVA.
@@
Wenhao Shi1thanks: Equal Contribution., Zhiqiang Hu2$\ast$, Yi Bin3,4, Junhua Liu2, Yang Yang1 See-Kiong Ng4, Lidong Bing, Roy Ka-Wei Lee2 1University of Electronic Science and Technology of China 2Singapore University of Technology and Design 3Tongji University 4National University of Singapore
@@ 1 Introduction
Motivation. Large language models (LLMs) exhibit impressive reasoning capabilities, drawing significant research interest in mathematical problem-solving in textual form Wei et al. (2022); Wang et al. (2023); Bin et al. (2023); Luo et al. (2023); Yue et al. (2023b); Gou et al. (2023); Zhou et al. (2023). However, the task of multimodal mathematical reasoning Lu et al. (2023) requires models to interpret diverse images and apply advanced reasoning skills. While open-source multimodal large language models (MLLMs) like LLaVA Liu et al. (2023) and Mini-GPT4 Zhu et al. (2023) perform well on visual question answering tasks Guo et al. (2023), they fall short of proprietary MLLMs (OpenAI, ; Google, ) in solving complex mathematical problems involving visual content.
Two common approaches to enhance MLLMs’ mathematical reasoning skills are prompt methods and fine-tuning methods. Prompt methods Lu et al. (2023); Wang et al. (2024b) leverage MLLMs’ latent abilities through carefully designed prompts, while fine-tuning methods Wang et al. (2024a); Hu et al. (2023a); Zheng et al. (2023) adjust model parameters using reasoning data collected from real-world or synthetic data from advanced LLMs (e.g., GPT-4). However, existing open-source image instruction fine-tuning datasets Lu et al. (2022b); Li et al. (2023); Lu et al. (2022a), which contain limited question-answer pairs per image, do not fully exploit visual information to enhance MLLMs’ multimodal mathematical reasoning capabilities.
@@
Contributions. Using the selected 40K data, the fine-tuned LLaVA-1.5 model, named Math-LLaVA-DS, achieved a significant improvement of 10.6% on MathVista Lu et al. (2023). To further enhance multimodal mathematical reasoning capabilities, we synthesized an additional 320K question-answer pairs based on the 40K images and seed questions, resulting in the MathV360K dataset. This comprehensive dataset, containing around 40K images and 360K question-answer pairs, significantly expands the coverage of multimodal mathematical reasoning. Fine-tuning LLaVA-1.5 with MathV360K, we developed Math-LLaVA, which outperforms the original LLaVA-1.5 by 19% on MathVista’s minitest split. We also evaluated Math-LLaVA on MMMU Yue et al. (2023a), demonstrating its improved generalizability.
@@ 2.1 Multimodal Large Language Models
The advancement of LLMs has spurred significant research interest in vision-language interaction, particularly in integrating visual knowledge into LLMs. The CLIP series Radford et al. (2021); Li et al. (2022) aligned visual and language modalities using contrastive learning on extensive image-text pairs. Recent studies increasingly use pre-training alignment and visual instruction tuning on LLMs for complex tasks like visual question answering and multimodal reasoning. MiniGPT-4 Zhu et al. (2023) engages in image-text dialogues by aligning visual features with text. Similarly, models like LLaVA Liu et al. (2023) and InstructBLIP Dai et al. (2024) use learnable projectors or query embeddings to interact with visual features. These approaches aimed to leverage high-quality pre-training and fine-tuning data to comprehend complex instructions. Models like mPLUG-Owl Ye et al. (2023), SPHINX Lin et al. (2023b), and MiniCPM-V2 Hu et al. (2024) introduced new grounding data types and modularization training to minimize hallucinations and enhance grounding abilities. Despite these advancements, MLLMs face challenges in solving multimodal mathematical problems using diagrams. Further exploration of the quality and format of image instructions is needed to improve the reasoning capabilities of MLLMs.
@@ 2.2 Multimodal Reasoning
The rapid development of MLLMs has advanced research on multimodal reasoning Chen et al. (2024a); You et al. (2023). Augmenting the original question and answer text data in restricted domains to further fine-tune MLLMs is a popular approach. For raw answers, rationales were either generated by humans Zhang et al. (2023c) or gathered from prominent LLMs Wang et al. (2024a); Lin et al. (2023a); Chen and Feng (2023). Additionally, VPD Hu et al. (2023b) proposed expanding answers by converting programming code formats to natural language formats. For raw questions, DDCoT Zheng et al. (2023) used LLMs to decompose the original questions into sub-questions. These methods, however, only utilize LLMs to target text-only data within restricted domains, neglecting to fully exploit the visual information in raw images for further enhancement. To evaluate the multimodal reasoning abilities of MLLMs more comprehensively, MathVista Lu et al. (2023), which involves various types of mathematical reasoning and skills, and MMMU Yue et al. (2023a), which encompasses multidisciplinary tasks, have been proposed. There is still significant room for improvement in open-source MLLMs.
@@ 4.2 Evaluation and Metrics
We evaluate our model using the minitest subset of MathVista Lu et al. (2023) in a zero-shot manner. This minitest subset comprises 1,000 samples, including 540 multiple-choice questions and 460 questions that require free-form answers in the form of integers, floats, or lists. MathVista adequately assesses the MLLMs’ multimodal mathematical skills, including algebraic reasoning (ALG), arithmetic reasoning (ARI), geometry reasoning (GEO), logical reasoning (LOG), numeric commonsense (NUM), scientific reasoning (SCI), and statistical reasoning (STA). Furthermore, MathVista questions can be categorized into the following subsets: FQA, GPS, MWP, TQA, and VQA. For evaluation, we first employ GPT-4 to extract the predicted choices or answers from responses, then report the answer accuracy, which entails determining whether the final answer matches the ground truth. Additionally, we evaluate our model’s enhanced generalizability using the MMMU benchmark Yue et al. (2023a). The MMMU benchmark, with 900 evaluation samples, encompasses six core disciplines: Art & Design, Business, Science, Health & Medicine, Humanities & Social Science, and Technology & Engineering, making it suitable for assessing the generalization of MLLMs’ reasoning capabilities.
@@ 5.1 Main Comparison on MathVista
We compare Math-LLaVA with other MLLMs on the minitest split of the MathVista benchmark in Table 1. As shown in the table, open-source MLLMs such as miniGPT4 Zhu et al. (2023), instructBLIP Dai et al. (2024), and LLaVA-1.5-13B have poor performance in multimodal mathematics, with overall accuracy lower than 30%. Compared to the base model, LLaVA-1.5-13B, with poor multimodal mathematical ability, Math-LLaVA achieves 46.6% overall accuracy with a significant improvement of 19%. More surprisingly, the proposed Math-LLaVA model outperforms close-source models Gemini 1.0 Pro Team et al. (2023) and Claude 3 Haiku Anthropic (2024), even achieving comparable performance to GPT-4V OpenAI , the most powerful close-source MLLMs. Interestingly, Math-LLaVA achieves 57.7% accuracy on GPS subset, outperforming G-LLaVA-13B Gao et al. (2023), which has been trained on 170K high-quality geometric image-caption and question-answer pairs. The superior performance of Math-LLaVA indicates that the data selection and synthesis of high-quality, diverse multimodal question-answer pairs are effective in improving MLLM’s multimodal mathematical reasoning capabilities.
@@ 5.2 Generalizability of Math-LLaVA
5.3 Overfitting to Synthesized Dataset
@@
Interestingly, we also observe that with text-only input, LLaVA-1.5-13B achieves an accuracy of 23.3% on MathVista. Potential reasons for this, as explored in Chen et al. (2024b), could be that visual content is unnecessary for many samples in MathVista and that unintentional data leakage may occur during the pre-training of LLMs and MLLMs.
@@ 6 Conclusions
We addressed the shortage of high-quality and diverse multimodal mathematical training datasets by creating MathV360K, which consists of 40K high-quality multimodal questions and answers from 24 existing datasets, along with 320K newly synthesized question-answer pairs. This comprehensive dataset enhances both the breadth and depth of multimodal mathematical questions. Using MathV360K, we fine-tuned Math-LLaVA, significantly improving its capability in multimodal mathematical reasoning, outperforming LLaVA-1.5 by 19 points on the minitest split of MathVista. Additionally, Math-LLaVA was validated on the MMMU benchmark, demonstrating its generalizability. Our research underscores the importance of dataset diversity and synthesis in enhancing the mathematical reasoning abilities of MLLMs.
@@ A.1 Source Data Statistics
A.2 Distribution Proportioning of Image Comprehension Complexity
@@ A.3 Cases Study
Additionally, we present several examples of newly generated questions, created by thoroughly mining images and questions from the selected dataset. As depicted in Figure 8, the existing dataset contains a limited number of image-question pairs. By fully utilizing the visual information from the images, we are able to generate a wider variety of questions from different perspectives, thereby enhancing the diversity of the problem set. The newly generated questions are created in a few-shot manner, referencing the format of existing question types. Consequently, these questions encompass more than just isolated visual content; they involve reasoning with the images. Moreover, the inclusion of complex questions, logically consistent rephrased questions, and simplified, underspecified questions increases the diversity and robustness of the dataset in terms of both question depth and format, compared to the original set of questions.

### v3

@@ Abstract
Large language models (LLMs) have demonstrated impressive reasoning capabilities, particularly in textual mathematical problem-solving. However, existing open-source image instruction fine-tuning datasets, containing limited question-answer pairs per image, do not fully exploit visual information to enhance the multimodal mathematical reasoning capabilities of Multimodal LLMs (MLLMs). To bridge this gap, we address the lack of high-quality, diverse multimodal mathematical datasets by collecting 40K high-quality images with question-answer pairs from 24 existing datasets and synthesizing 320K new pairs, creating the MathV360K dataset, which enhances both the breadth and depth of multimodal mathematical questions. We introduce Math-LLaVA, a LLaVA-1.5-based model fine-tuned with MathV360K. This novel approach significantly improves the multimodal mathematical reasoning capabilities of LLaVA-1.5, achieving a 19-point increase and comparable performance to GPT-4V on MathVista’s minitest split, and yielding leading performance on Math-V and MathVerse. Furthermore, Math-LLaVA demonstrates enhanced generalizability, showing substantial improvements on the MMMU benchmark. Our research highlights the importance of dataset diversity and synthesis in advancing MLLMs’ mathematical reasoning abilities. The code and data are available at: https://github.com/HZQ950419/Math-LLaVA.
@@
Wenhao Shi1thanks: Equal Contribution., Zhiqiang Hu2$\ast$, Yi Bin3,4thanks: The corresponding author, email: yi.bin@hotmail.com, Junhua Liu2, Yang Yang1 See-Kiong Ng4, Lidong Bing, Roy Ka-Wei Lee2 1University of Electronic Science and Technology of China 2Singapore University of Technology and Design 3Tongji University 4National University of Singapore
@@ 1 Introduction
Motivation. Large language models (LLMs) exhibit impressive reasoning capabilities, drawing significant research interest in mathematical problem-solving in textual form Zhang et al. (2020); Wei et al. (2022); Wang et al. (2023); Bin et al. (2023); Luo et al. (2023); Yue et al. (2023b); Gou et al. (2023); Zhou et al. (2023). However, the task of multimodal mathematical reasoning Lu et al. (2023) requires models to interpret diverse images and apply advanced reasoning skills. While open-source multimodal large language models (MLLMs) like LLaVA Liu et al. (2023) and Mini-GPT4 Zhu et al. (2023) perform well on VQA tasks Guo et al. (2023), they fall short of proprietary MLLMs (OpenAI, ; Google, ) in solving complex mathematical problems involving visual content.
Two common approaches to enhance MLLMs’ mathematical reasoning skills are prompt methods and fine-tuning methods. Prompt methods Lu et al. (2023); Wang et al. (2024c) leverage MLLMs’ latent abilities through carefully designed prompts, while fine-tuning methods Wang et al. (2024b); Hu et al. (2023a); Zheng et al. (2023) adjust model parameters using reasoning data collected from real-world or synthetic data from advanced LLMs (e.g., GPT-4). However, existing open-source image instruction fine-tuning datasets Lu et al. (2022b); Li et al. (2023); Lu et al. (2022a), which contain limited question-answer pairs per image, do not fully exploit visual information to enhance MLLMs’ multimodal mathematical reasoning capabilities.
@@
Contributions. Using the selected 40K data, the fine-tuned LLaVA-1.5 model, named Math-LLaVA-DS, achieved a significant improvement of 10.6% on MathVista Lu et al. (2023). To further enhance multimodal mathematical reasoning capabilities, we synthesized an additional 320K question-answer pairs based on the 40K images and seed questions, resulting in the MathV360K dataset. This comprehensive dataset, containing around 40K images and 360K question-answer pairs, significantly expands the coverage of multimodal mathematical reasoning. Fine-tuning LLaVA-1.5 with MathV360K, we developed Math-LLaVA, which outperforms the original LLaVA-1.5 by 19% on MathVista’s minitest split and achieves leading performance on Math-V Wang et al. (2024a) and MathVerse Zhang et al. (2024). We also evaluated Math-LLaVA on MMMU Yue et al. (2023a), demonstrating its improved generalizability.
@@ 2.1 Multimodal Large Language Models
The advancement of LLMs has spurred significant research interest in vision-language interaction, particularly in integrating visual knowledge into LLMs. The CLIP series Radford et al. (2021); Li et al. (2022) aligned visual and language modalities using contrastive learning on extensive image-text pairs. Recent studies increasingly use pre-training alignment and visual instruction tuning on LLMs for complex tasks like visual question answering, artwork analysis, and multimodal reasoning Li et al. (2024a); Bin et al. (2024). MiniGPT-4 Zhu et al. (2023) engages in image-text dialogues by aligning visual features with text. Similarly, models like LLaVA Liu et al. (2023) and InstructBLIP Dai et al. (2024) use learnable projectors or query embeddings to interact with visual features. These approaches aimed to leverage high-quality pre-training and fine-tuning data to comprehend complex instructions. Models like mPLUG-Owl Ye et al. (2023), SPHINX Lin et al. (2023b), and MiniCPM-V2 Hu et al. (2024) introduced new grounding data types and modularization training to minimize hallucinations and enhance grounding abilities. Despite these advancements, MLLMs face challenges in solving multimodal mathematical problems using diagrams. Further exploration of the quality and format of image instructions is needed to improve the reasoning capabilities of MLLMs.
@@ 2.2 Multimodal Reasoning
The rapid development of MLLMs has advanced research on multimodal reasoning Chen et al. (2024a); You et al. (2023). Augmenting the original question and answer text data in restricted domains to further fine-tune MLLMs is a popular approach. For raw answers, rationales were either generated by humans Zhang et al. (2023c) or gathered from prominent LLMs Wang et al. (2024b); Lin et al. (2023a); Chen and Feng (2023); Li et al. (2024b). Additionally, VPD Hu et al. (2023b) proposed expanding answers by converting programming code formats to natural language formats. For raw questions, DDCoT Zheng et al. (2023) used LLMs to decompose the original questions into sub-questions. These methods, however, only utilize LLMs to target text-only data within restricted domains, neglecting to fully exploit the visual information in raw images for further enhancement. To evaluate the multimodal reasoning abilities of MLLMs more comprehensively, MathVista Lu et al. (2023), Math-V Wang et al. (2024a) and MathVerse Zhang et al. (2024), which involve various types of mathematical reasoning and skills, and MMMU Yue et al. (2023a), which encompasses multidisciplinary tasks, have been proposed. There is still significant room for improvement in open-source MLLMs.
@@ 4.2 Evaluation and Metrics
We evaluate our model using the minitest subset of MathVista Lu et al. (2023) in a zero-shot manner. This minitest subset comprises 1,000 samples, including 540 multiple-choice questions and 460 questions that require free-form answers in the form of integers, floats, or lists. MathVista adequately assesses the MLLMs’ multimodal mathematical skills, including algebraic reasoning (ALG), arithmetic reasoning (ARI), geometry reasoning (GEO), logical reasoning (LOG), numeric commonsense (NUM), scientific reasoning (SCI), and statistical reasoning (STA). Furthermore, MathVista questions can be categorized into the following subsets: FQA, GPS, MWP, TQA, and VQA. For evaluation, we first employ GPT-4 to extract the predicted choices or answers from responses, then report the answer accuracy, which entails determining whether the final answer matches the ground truth. We also conduct evaluation using the Math-V Wang et al. (2024a) and MathVerse Zhang et al. (2024). Math-V is a meticulously curated collection of 3,040 mathematical problems with visual contexts sourced from real math competitions. MathVerse consists of 2,612 multi-subject math problems varying degrees of information content in multi-modality. Additionally, we evaluate our model’s enhanced generalizability using the MMMU benchmark Yue et al. (2023a). The MMMU benchmark, with 900 evaluation samples, encompasses six core disciplines: Art & Design, Business, Science, Health & Medicine, Humanities & Social Science, and Technology & Engineering, making it suitable for assessing the generalization of MLLMs’ reasoning capabilities.
@@ 5.1 Main Comparison on MathVista
We compare Math-LLaVA with other MLLMs on the minitest split of the MathVista benchmark in Table 1. As shown in the table, open-source MLLMs such as miniGPT4 Zhu et al. (2023), instructBLIP Dai et al. (2024), and LLaVA-1.5-13B have poor performance in multimodal mathematics, with overall accuracy lower than 30%. Compared to the base model, LLaVA-1.5-13B, with poor multimodal mathematical ability, Math-LLaVA achieves 46.6% overall accuracy with a significant improvement of 19%. More surprisingly, the proposed Math-LLaVA model outperforms close-source models Gemini 1.0 Pro Team et al. (2023) and Claude 3 Haiku Anthropic (2024), even achieving comparable performance to GPT-4V OpenAI , the most powerful close-source MLLMs. Interestingly, Math-LLaVA achieves 57.7% accuracy on GPS subset, outperforming G-LLaVA-13B Gao et al. (2023a), which has been trained on 170K high-quality geometric image-caption and question-answer pairs. The results on Math-V are shown in Table 2. Math-LLaVA also achieves significant improvement compared to the base model and leading performance than Qwen-VL-Max Bai et al. (2023) and most open-source MLLMs. The results on MathVerse could be found from Table 7 in Appendix. The superior performance of Math-LLaVA indicates that the data selection and synthesis of high-quality, diverse multimodal question-answer pairs are effective in improving MLLM’s multimodal mathematical reasoning capabilities.
@@
Model Math-V
ALL Alg AnaG Ari CG Comb Cnt DG GT Log Angle Area Len SG Sta Topo TG
Heuristics Baselines
Random Chance 7.2 1.5 11.9 7.1 9.7 4.8 6.0 22.1 1.1 7.6 0.6 9.4 6.7 8.2 8.6 13.0 7.1
Human 68.8 55.1 78.6 99.6 98.4 43.5 98.5 91.3 62.2 61.3 33.5 47.2 73.5 87.3 93.1 99.8 69.0
Close-Source Multimodal Large Langugae Models (MLLMs)
Qwen-VL-Plus Bai et al. (2023) 10.7 11.3 17.9 14.3 12.7 4.8 10.5 15.4 8.9 14.3 11.6 6.4 10.0 14.3 6.9 8.7 11.3
Qwen-VL-Max Bai et al. (2023) 15.6 10.7 19.1 20.0 16.9 12.5 17.9 16.4 12.2 21.0 13.3 14.2 19.8 11.5 20.7 13.0 17.3
Gemini Pro Team et al. (2023) 17.7 15.1 10.7 20.7 20.1 11.9 7.5 20.2 21.1 16.8 19.1 19.0 20.0 14.3 13.8 17.4 20.8
GPT-4V OpenAI 22.8 27.3 32.1 35.7 21.1 16.7 13.4 22.1 14.4 16.8 22.0 22.2 20.9 23.8 24.1 21.7 25.6
Open-Source Multimodal Large Langugae Models (MLLMs)
SPHINX-V2-13B Lin et al. (2023b) 9.7 6.7 7.1 12.9 7.5 7.7 6.0 9.6 16.7 10.1 11.0 11.8 12.5 8.2 8.6 8.7 6.0
LLaVA-1.5-13B Liu et al. (2024) 11.1 7.0 14.3 14.3 9.1 6.6 6.0 13.5 5.6 13.5 10.4 12.6 14.7 11.5 13.8 13.0 10.7
Math-LLaVA 15.7 9.0 20.2 15.7 18.2 10.1 10.5 16.4 14.4 16.0 20.2 18.4 17.6 9.4 24.1 21.7 17.9
Table 2: Performance Comparison on the Math-V benchmark with the accuracy metric across various mathmatical subjects. Baseline results are obtained from Wang et al. (2024a). The best results in both the close-source and open-source MLLMs are in bold.
@@ 5.2 Generalizability of Math-LLaVA
5.3 Overfitting to Text Modality
@@
Interestingly, we also observe that with text-only input, LLaVA-1.5-13B achieves an accuracy of 23.3% on MathVista. Potential reasons as explored in Chen et al. (2024b) could be that visual content is unnecessary for many samples in MathVista and that unintentional data leakage may occur during the pre-training of LLMs and MLLMs.
@@ 6 Conclusions
We addressed the shortage of high-quality and diverse multimodal mathematical training datasets by creating MathV360K, which consists of 40K high-quality multimodal questions and answers from 24 existing datasets, along with 320K newly synthesized question-answer pairs. This comprehensive dataset enhances both the breadth and depth of multimodal mathematical questions. Using MathV360K, we fine-tuned Math-LLaVA, significantly improving its capability in multimodal mathematical reasoning, outperforming LLaVA-1.5 by 19 points on the minitest split of MathVista, and yielding leading performance on Math-V and MathVerse. Additionally, Math-LLaVA was validated on the MMMU benchmark, demonstrating its generalizability. Our research underscores the importance of dataset diversity and synthesis in enhancing the mathematical reasoning abilities of MLLMs.
@@ 8 Ethics Statement
9 Acknowledgement
This work is supported by the National Natural Science Foundation of China under grant 62220106008, U20B2063, and 62102070. This research is supported by A*STAR, CISCO Systems (USA) Pte. Ltd and National University of Singapore under its Cisco-NUS Accelerated Digital Economy Corporate Laboratory (Award I21001E0002).
@@ A.1 Source Data Statistics
A.2 Results on MathVerse Benchmark
The proposed Math-LLaVA model has demonstrated impressive performance on MathVista and Math-V. To assess its multimodal mathematical reasoning capabilities more comprehensively, we conduct evaluation experiments using the MathVerse benchmark Zhang et al. (2024). The results are shown in Table 7. Math-LLaVA also achieves significant improvement compared to the base model and impressive performance among most open-source MLLMs.
Model MathVerse
ALL Text Dominant Text Lite Vision Intensive Vision Dominant Vision Only
Heuristics Baselines
Random Chance 12.4 12.4 12.4 12.4 12.4 12.4
Human 64.9 71.2 70.9 61.4 68.3 66.7
Close-Source Multimodal Large Langugae Models (MLLMs)
Qwen-VL-Plus Bai et al. (2023) 11.8 15.7 11.1 9.0 13.0 10.0
Gemini Pro Team et al. (2023) 23.7 27.6 23.7 19.4 20.3 20.5
GPT-4V OpenAI 38.9 52.1 40.9 34.9 33.6 29.8
Open-Source Multimodal Large Langugae Models (MLLMs)
mPLUG-Owl-7B Ye et al. (2023) 4.6 6.6 6.3 6.3 5.6 4.9
LLaMA-Adapter-V2-7B Gao et al. (2023b) 5.7 6.2 5.9 6.1 4.2 6.1
LLaVA-1.5-13B Liu et al. (2024) 7.6 8.8 7.6 7.4 7.4 6.9
SPHINX-V2-13B Lin et al. (2023b) 12.2 13.9 11.6 11.6 13.5 10.4
Math-LLaVA 19.8 22.3 21.6 20.8 19.2 15.2
Table 7: Performance Comparison on the MathVerse benchmark with the accuracy metric. Baseline results are obtained from Zhang et al. (2024). The best results in both the close-source and open-source MLLMs are in bold.
A.3 Distribution Proportioning of Image Comprehension Complexity
@@ A.4 Cases Study
Additionally, we present several examples of newly generated questions, created by thoroughly mining images and questions from the selected dataset. As depicted in Figure 8, existing dataset contains a limited number of image-question pairs. By fully utilizing the visual information from the images, we are able to generate a wider variety of questions from different perspectives, thereby enhancing the diversity of the problem set. The generated questions are created in a few-shot manner, referencing the format of existing question types. Consequently, these questions encompass more than just isolated visual content; they involve reasoning with the images. Moreover, the inclusion of complex questions, logically consistent rephrased questions, and simplified, underspecified questions increases the diversity and robustness of the dataset in terms of both question depth and format, compared to the original set of questions.